how do you solve minimize p=-21x+11y for the constraints from excercise 9.
Exercise 9 : x is greater than or equal to 0
y is greater than or eaual to 0
y is greater than or equal to 4x-4
y is less-than or equal to x+5

Answer by robertb(5830)

The feasibility region is all the points inside and on the boundary of the quadrilateral having vertices or corner points (0,0), (1,0), (3,8), (0,5). By the fundamental theorem of linear programming the maximum and minimum will occur at the corner points.
At (0,0), p = 0.
At (1,0), p = -21
At (3,8), p = -21*3 + 11*8 = -63 + 88 = 25
At (0,5), p= 55.
Thus the maximum of p is 55 and the minimum is -21.